Calculate the number of moles and the mass of negative ions present in 15.67 g of magnesium nitrate, Mg(NO3)2.
Answer #1

molar mass of Mg(NO3)2 = 148.3 g/mol

no. of moles = 15.67/148.3 = 0.105 moles

One mole of Mg(NO3)2 ---> Mg 2+ + 2 NO3- conrtains one mole of Mg and 2 mol of NO3-

no.of moles of Anion = 0.105*2 = 0.21 mol

mass of NO3 2- = 62+2 =64

weight -= 0.21* 64 = 13.44 gram